OpenStudy (thomas5267):

Prove the following \[ \lim_{x\to 0^+}x\ln(x)=0 \]

OpenStudy (thomas5267):

The simple brute force application of L'Hoptial rule won't work \[ x\ln(x)=\frac{x}{\left(\ln(x)\right)^{-1}}\\ \lim_{x\to0^+}\frac{x}{\left(\ln(x)\right)^{-1}}=\lim_{x\to0^+}\frac{-1}{x\left(\ln(x)\right)^2} \]

OpenStudy (freckles):

\[\lim_{x \rightarrow 0^+} \frac{\ln(x)}{\frac{1}{x}}=\lim_{x \rightarrow 0^+}\frac{\frac{1}{x}}{\frac{-1}{x^2}}=\lim_{x \rightarrow 0^+} -x=0\]

OpenStudy (thomas5267):

Really clever!

OpenStudy (freckles):

Is that really proving the limit is 0 though? Or is it just finding the limit?

OpenStudy (irishboy123):

what's the difference?

OpenStudy (freckles):

for example: \[\lim_{x \rightarrow 2} \frac{x^2-4}{x-2}=4 \text{ we can find this limit algebraically } \\ \text{ but this isn't called a proof }\]

OpenStudy (freckles):

we can also find the limit using l'hospital too but I'm also thinking that is not really considered a proof

OpenStudy (irishboy123):

sadly, it is

OpenStudy (irishboy123):

so i think

OpenStudy (freckles):

why do we need to prove the limit if we find it algebraically but we finding the limit using l'hospital means we have done the proof already..

OpenStudy (freckles):

it looks like someone try to use squeeze theorem

OpenStudy (freckles):

the one with 5 votes

OpenStudy (irishboy123):

which makes sense to me historically, limits came from trying to make sense of calculus, ie dividing by zero, which is what calculus does so l'Hopital is calculus proving limits. squeeze is geometry.

OpenStudy (irishboy123):

the one at the bottom is what occurred to me but i don't have the math.

OpenStudy (freckles):

I really like that squeeze theorem junk he did

OpenStudy (freckles):

Maybe if we look harder we can do this for all limits that exist

OpenStudy (freckles):

you know apply squeeze theorem I know sometimes it might be tough

OpenStudy (irishboy123):

but he's integrating, so he's using limits to prove a limit 😰

OpenStudy (freckles):

I must eat I might return

OpenStudy (irishboy123):

the sin x / x limit that needs squeeze that is the best example IMHO of where the circularity occurs.

OpenStudy (thomas5267):

I am asking a proof. \(\lim_{x\to0+}x\ln(x)=0\) is a statement which could be true or false. Therefore, you have to prove it. To express it as a find statement, I would have wrote the following Find the following limit. \[ \lim_{x\to0+}x\ln(x) \] It is pretty much a pointless distinction.

OpenStudy (thomas5267):

I thought squeeze theorem could be proved independently of the limit you are evaluating. I would not be surprised if you could derive it directly from the delta-epsilon definition of limits. @IrishBoy123

OpenStudy (thomas5267):

What is wrong with the integration and limit answer? He simply transformed a limit to a easier limit, just like we did using L'Hopital. Where is the circular reasoning? I thought l'Hopital is calculus proving limit but then I checked the proof on wiki. Does not seem like this is the case. https://en.wikipedia.org/wiki/L'H%C3%B4pital's_rule

OpenStudy (thomas5267):

For completeness, https://en.wikipedia.org/wiki/Squeeze_theorem Squeeze theorem can be proved using delta-epsilon alone.